OpenStudy (anonymous):

Complete the Next Step: A partial solution to the equation 4d − 12 = 12 − 4d is given below. Choose the step that is missing from the solution. 4d − 12 = 12 − 4d 4d − 24 = −4d ? 3 = d

OpenStudy (anonymous):

A. 24 = −8d B. −24 = 8d C .−24 = 0d D. −24 = −8d

OpenStudy (anonymous):

D is the answer

OpenStudy (anonymous):

D. -24 = -8d

OpenStudy (anonymous):

but that could be simplified too right?

OpenStudy (anonymous):

4d − 12 = 12 − 4d 4d+4d-12=12 8d=12+12 8d=24 d=24/8 d=3

OpenStudy (anonymous):

It ask for the next step

OpenStudy (anonymous):

every step is done at there....